OpenStudy (anonymous):

From 1940 to 1980 the consumer price index (CPI) increased from 28.6 to 89.4. If a gallon of milk cost $0.75 in 1940 and the price of milk increased at the same rate as the CPI from 1940 to 1980, approximately how much did a gallon of milk cost in 1980? I need help getting this started. can someone help me? :)

OpenStudy (mertsj):

Write a proportion.

OpenStudy (mertsj):

\[\frac{28.6}{89.4}=\frac{.75}{x}\]

OpenStudy (anonymous):

I may be wrong, but it looks like it's asking for the rate in change over the past forty years. Then multiply that by the cost of milk to arive at the adjusted price in 1980

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (anonymous):

Thank you, :)

OpenStudy (anonymous):

well, I don't think this is a division problem, more of a subtration

OpenStudy (anonymous):

1980-1940 = 40 years 89.4 - 28.6 = 60.80

OpenStudy (anonymous):

now you could do your division 60.80/40= 1.52 and multiply that by the cost of milk in 1940 0.75 * 1.52 =$1.14

OpenStudy (anonymous):

Thank you! I was doing that but somewhere I got the wrong answer! Thank you! :)

OpenStudy (anonymous):

and if that is not the answer then I might think to add the 1.52 to the 0.75 =$2.27